a spherical liquid drop has diameter of 2cm and is given a charge of 1 micro coulomb. if 2 such drops coalesce to form a single drop. what is the potential at the surface of the drop so formed.

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
.
Lets Give the charge on every little drop a chance to be q C. 
In this way, VsmallVsmall=kqr→q=Vrk=kqr→q=Vrk 
where VsmallVsmall is the potential on every little circle. 
Add up to volume of every single little circle = Volume of vast drop 
∴∴27×4/3πr3=4/3πR^327×4/3πr3=4/3πR^3 
∴3r=R∴3r=R 
∴Vlarge=kQ/R=k27q/3r=9Vsmall
.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)